please help me
solve this algebra equation jean and chinoso's age add up to 29, 7 years ago jean was as old as chinoso find their present ages ?

Answer by math_helper(2461)

(1) J = C + 7 (from "7 years ago Jean was as old as Chinoso")
(2) J + C = 29 (…"Jean and Chinoso's age add up to 29")

Substitute for J:
(C+7) + C = 29
2C = 22 —> C = 11 —> J = 18
———
Ans: Chinoso is 11 and Jean is 18
———
Check: 11+18 = 29
========================

Alternate (shorter) solution, using Jean's age (expressed in terms of C):
C+7 = (29-C) —> 2C = 22 —> C=11 —> J=18

…and using Chinoso's age (expressed in terms of J):
J-7 = 29-J —> 2J = 36 —> J=18 —> C=11